OpenStudy (anonymous):

Suppose that A>0 is a real number and that B = lnA. Find k so that 3^B = A^k......Please help :(

OpenStudy (anonymous):

Can use logarithm to solve for k in terms of A.

hartnn (hartnn):

taking ln on both sides B ln 3 = k ln A but B= ln A ln A ln 3= k ln A so k =ln 3 got this?

OpenStudy (anonymous):

How did you get the B out from the exponent, I got it to ln3^B = lnA^k but from there was confused.

hartnn (hartnn):

ok property of logarithms is : \(ln \: a^b=b \: ln \:a\) ok?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

Ha, I did it the long way. I did it using log_{A}, but still got the same answer.

OpenStudy (pradipgr817):

how B came out hartnn

hartnn (hartnn):

u mean in the property? @pradipgr817